Kiera wants to make a solution that is 60% Alcohol. How many liters of a 70% Alcohol solution should Kiera add to 12 liters of a 45% Alcohol solution to achieve her goal?
Answer by mananth(16946)

---------------- percent ---------------- quantity
Alcohol I 45 ---------------- 12 liters
Alcohol II 70 ---------------- x liters
Mixture 60 ---------------- 12 + x liters

45 * 12 + 70 x = 60 ( 12 + x )
540 + 70 x = 720 + 60 x
70 x + -60 x = 720 - -540
10 x = 180
/ 10
x = 18
18 liters of 70 % Alcohol